The perimeter of a rectangle is 68cm. If its breadth is 4cm less than its length ,find the length and breadth

Answer by solve_for_x(190)

Let L represent the length.

The breadth is then L - 4.

The perimeter is equal to twice the length, plus twice the breadth:

2L + 2(L - 4) = 68

2L + 2L - 8 = 68

4L - 8 = 68

4L = 76

L = 19

Then, the breadth is L - 4 = 19 - 4 = 15

Solution: Length = 19, Breadth = 15

Check: 19 + 15 + 19 + 15 = 68